# Gamma positivity, PL homeomorphism types, and orthogonal polynomials

### 23 Mar 2026 v2

@@ Introduction
In order to have a better grasp of how we can think about these objects objects quantitatively, we considered averaging and asymptotic properties. This does not seem to have been studied in many settings in the context of gamma vectors apart from work on symmetric edge polytopes in [15]. More specifically, we focused on measuring how important the flag property is to nonnegativity of gamma vectors of simplicial spheres. To this end, we replaced the flag property by the more general link condition $\operatorname{lk}_{\Delta}(a)\cap\operatorname{lk}_{\Delta}(b)=\operatorname{lk}_{\Delta}(e)$ for edges $e\in\Delta$ of a simplicial complex $\Delta$. This is a common combinatorial application of the flag property which ends up being equivalent to preservation of piecewise linear (PL) homeomorphism types of compact PL manifolds without boundary [37].
Given a PL simplicial sphere, the characterization of possible $g$-vectors of simplicial spheres (Theorem 1.6) and approximations of pseudopowers arising from Macaulay representations (Definition 1.8) discussed in Section 2.1 (e.g. applications of results in [17] and [10]) enable us to study the effect of the link condition on the structure of simplicial spheres with nonnegative gamma vectors in Section 2.2. The results involved also apply to generalizations of gamma positivity coming from connections between orthogonal polynomials and lattice paths (see Section 3). This fits into lattice path-related properties of gamma vector transformations (Proposition 2.3), recent gamma positivity-related work (e.g. [11] and [16]), and positivity and structural results of the gamma polynomial from inverted Chebyshev expansions (Theorem 1.2, Corollary 1.35 on p. 25 of [41] and Corollary 1.10 on p. 9 of [44]). The main results are listed below.
@@
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial sphere satisfying the link condition such that $\gamma(\Delta)\geq 0$. Given an edge $e\in\Delta$, let $\widetilde{\Delta}_{e}$ be the contraction of $e$.
@@
    For gamma vectors $\gamma(\Delta)$, lower bounds on $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ can be phrased as lower bounds on $\gamma_{k+2}(\Delta)$ since this term only appears in $g_{k+2}(\Delta)$ (see Proposition 2.1). However, the asymptotic effect on high-dimensional spheres satisfying the link condition with $\gamma(\Delta)\geq 0$ is the largest when the gamma vector components involved are asymptotically dominant in the $g$-vector components of the same index. For example, this occurs if $\gamma_{\ell+1}\geq(d+1)g_{\ell}(\Delta)$ (e.g. when $g_{\ell+1}(\Delta)\geq 2(d+1)g_{\ell}(\Delta)$). While the results discussed here involve individual inequalities associated to constant indices $k$, the underlying tools (Proposition 2.40, Corollary 2.42, Corollary 2.45) apply to the behavior of collections of inequalities associated to varying indices.
@@ 1 Background
      $g\coloneq(h_{0},h_{1}-h_{0},h_{2}-h_{1},\ldots,h_{\left\lfloor\frac{d}{2}\right\rfloor}-h_{\left\lfloor\frac{d}{2}\right\rfloor-1})$ satisfies the inequalities $0\leq g_{i+1}\leq g_{i}^{\langle i\rangle}$ for all $0\leq i\leq\left\lfloor\frac{d}{2}\right\rfloor-1$ (i.e. $g$ is the $h$-vector of some Cohen-Macaulay simplicial complex) (Theorem 2.2 on p. 56 of [46] and Theorem 1.10).
Remark 1.7.
(Piecewise linear (PL) assumption)
By “simplicial sphere”, we will mean a piecewise linear (PL) simplicial sphere (see p. 4 of [27] and p. 26 of [25]).
@@
Apart from realizations of simplicial spheres (Theorem 1.6), these expansions are involved in $f$-vectors and $h$-vectors attained by general classes of simplicial complexes.
@@ 2.1 General Cohen-Macaulay simplicial complexes and simplicial spheres
In this subsection, we will describe the construction of nonnegative gamma vectors realized by simplicial spheres and Cohen–Macaulay simplicial complexes in general and show that sufficiently large ones can be extended component by component (Corollary 2.13 and Remark 2.14). The key input comes from estimates of terms from Macaulay representations of $g$-vectors and $h$-vectors (Corollary 2.8) and the $g$-conjecture for simplicial spheres (Theorem 1.6). If $\gamma_{0}=1$ and $\gamma_{1},\ldots,\gamma_{\frac{d}{2}}\geq 0$ that grow sufficiently quickly relative to the dimension with $\gamma_{j}$ while satisfying designated order $\frac{j}{j-1}$ upper bounds in $\gamma_{1},\ldots,\gamma_{j-1}$, they give rise to nonnegative gamma vectors $(\gamma_{0},\gamma_{1},\ldots,\gamma_{\frac{d}{2}})$ realized by simplicial spheres.
@@
    The first inequality can be rewritten as
@@
    where p=d2sp=d-2s and q=rsq=r-s. Note that $p\geq 2q$ since $0\leq r\leq\frac{d}{2}$. Writing
@@
    $\gamma_{\ell+1}\geq b_{\ell+1,0}+b_{\ell+1,1}\gamma_{1}+\ldots+b_{\ell+1,\ell-1}\gamma_{\ell-1}+b_{\ell+1,\ell}\gamma_{\ell}.$
@@
While there is a lower bound $g_{1}(\Delta)\geq d-1$ for (d1)(d-1)-dimensional simplicial spheres satisfying the link condition (attained by the boundary of the $d$-dimensional cross polytope), the parameter $g_{1}(\Delta)$ can grow arbitrarily quickly with respect to $d$. In other words, $d$ can grow arbitarily slowly with respect to $g_{1}(\Delta)$ even when it is not a constant. We can see this by considering flag simplicial spheres, which must satisfy the link property. This is because flagness is preserved by edge subdivision and each edge subdivision increases the number of vertices by 1.
@@
    The $h_{i}(\Delta)$ statement follows from $h_{i}(\Delta)=1+g_{1}(\Delta)+\ldots+g_{i}(\Delta)$.
@@
    This follows from the same reasoning as Part 1 after replacing $g_{i}(\Delta)$ by $h_{i}(\Delta)$, using $a_{i,0}$ instead of $b_{i,0}$ in Proposition 2.1, and applying Theorem 1.10 instead of Theorem 1.6.
@@
If $\gamma_{1}=\cdots=\gamma_{i}=0$, we have equality. Suppose that $\gamma_{k}>0$ for some $1\leq k\leq i$. This would mean that both sides are strictly positive. Let $\alpha=b_{i,0}+b_{i,1}\gamma_{1}+\ldots+b_{i,i-1}\gamma_{i-1}+\gamma_{i}$ and
@@
Taking index changes into account, the $f$-vector realizability conditions of Theorem 1.9 state that we need $\gamma_{i+2}(\Delta)\leq\gamma_{i+1}(\Delta)^{\langle i+1\rangle}$. Substituting in i1i-1 in place of $i$, we have $\gamma_{i+1}(\Delta)\leq\gamma_{i}(\Delta)^{\langle i\rangle}$. The steps of the proof of Corollary 2.13 applying the estimates of Corollary 2.8 imply the controlled growth statement analogous to these results.
@@
If $\Delta$ is a simplicial sphere, Part 3 of Theorem 1.6 implies that $g(\Delta)=h(T)$ for some Cohen–Macaulay simplicial complex $T$. This implies the following:
@@
    This follows directly from Part 3 of Theorem 1.6.
@@ 2.2 Simplicial spheres satisfying the link condition
This is a natural condition to consider from the perspective of piecewise linear (PL) homeomorphisms since they can be decomposed into edge subdivisions and contractions (Corollary [10:2d] on p. 302 of [2], Corollary 4.1 on p. 75 of [31]) and the link condition is equivalent to the contraction of every edge preserving the PL homeomorphism class of a PL manifold without boundary [37]. In addition, the condition $\operatorname{lk}_{\Delta}(a)\cap\operatorname{lk}_{\Delta}(b)=\operatorname{lk}_{\Delta}(a,b)$ for edges $\{a,b\}\in\Delta$ is a common application of flagness, which ties many examples from gamma positivity together (see [4], [20]).
@@
In particular, this comes from a shift in perspective on these global/averaged inequalities as information on ratios of $g$-vector components $\frac{g_{\ell+1}(\Delta)}{g_{\ell}(\Delta)}$. Given this situation, we mainly focus on the global/averaged version of the $M$-vector condition on the $g$-vectors of the edge contractions $\widetilde{\Delta}_{e}$ of edges $e\in\Delta$. We note that individual local conditions induce “sandwiching” of local parts (Proposition 2.28) with global/averaged counterparts (Corollary 2.31) when combined with the PL manifold condition. We will primarily focus on the lower bounds and their comparison with nonnegativity of the $g$-vectors of the contractions $\widetilde{\Delta}_{e}$ of edges $e\in\Delta$ (Corollary 2.33) since this is the bound specific to spheres rather than PL manifolds in general.
The general shape of the global/averaged $M$-vector conditions on $g$-vectors of contractions $\widetilde{\Delta}_{e}$ of edges $e\in\Delta$ involves a lower bound on $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ in terms of $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$, the number of edges $f_{1}(\Delta)^{\frac{1}{k}}$, and $\frac{g_{k}(\Delta)^{\frac{k+1}{k}}}{g_{k+1}(\Delta)}$ (Proposition 2.38 and Example 2.39). The last parameter measures how far the $M$-vector condition $g_{k+1}(\Delta)\leq g_{k}(\Delta)^{\langle k\rangle}$ is form equality (Part 2c) of Theorem 1.6 and Corollary 2.8). Increasing these parameters decreases the resulting lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$.
@@ 2.2.1 Main underlying inequalities
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial complex. Given an edge $e\in\Delta$, let $\widetilde{\Delta}_{e}$ be the contraction of $e$ in $\Delta$. Then, we have
@@
    $\beta_{k}\geq\frac{2\left(2d(d-1)-\left(\frac{d}{2}-1\right)\left(\frac{d}{2}+2\right)\right)}{2}\approx\frac{7}{2}.$
@@
    Suppose that $\Delta$ is a simplicial sphere satisfying the link condition. In order for the “averaged” global version of the inequalities $g(\widetilde{\Delta}_{e})\geq 0$ for all the edges $e\in\Delta$ from Corollary 2.25 to impose a nontrivial condition on the $g_{k+2}(\Delta)$ beyond the $M$-vector condition on the $g$-vectors of arbitrary simplicial spheres (Part 2c) of Theorem 1.6), we need $g_{2}(\Delta)<(d^{2}-2)-(d-3)g_{1}(\Delta)$. By the $g(\Delta)\geq 0$ assumption, this also implies that $g_{1}(\Delta)<(d+3)+\frac{7}{d-3}$. If $d\geq 4$, the latter bound implies that $g_{1}(\Delta)<d+10$. This implies that $g_{2}(\Delta)\leq\binom{d+11}{2}$, which is a stronger bound for $d\gg 0$. Since there is a linear bound on $g_{1}(\Delta)$, each $g_{\ell}(\Delta)$ with $0\leq\ell\leq\frac{d}{2}$ is bounded above by a degree $\ell$ polynomial in $d$ (Corollary 2.11).
@@
    Since the coefficient of each $g$-vector component in Part 1 of Corollary 2.25 is strictly positive and $g(\Delta)\geq 0$, having $g_{k+1}(\Delta)=0$ would imply that $g_{k+2}(\Delta)=0$ and $g_{k}(\Delta)=0$. If $g_{\ell}(\Delta)=0$ for some $1\leq\ell\leq\frac{d}{2}$. this would mean that $g_{k}(\Delta)=0$ for all $0\leq k\leq\frac{d}{2}$. However, this is impossible in our setting.
@@
    We can compare these to initial bounds from the $M$-vector condition on $g$-vectors of arbitrary simplicial spheres. Since $g_{k+2}(\Delta)\leq g_{k+1}(\Delta)^{\langle k+1\rangle}\leq g_{k+1}(\Delta)^{\frac{k+2}{k+1}}$ and $g_{k+1}(\Delta)\leq g_{1}(\Delta)^{\frac{1}{k+1}}\Longrightarrow g_{k+1}(\Delta)^{\frac{1}{k+1}}\leq g_{1}(\Delta)$ by Corollary 2.8 and Corollary 2.9, we have
@@
    If $\gamma(\Delta)\geq 0$, then Lemma 2.3 implies that $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}\geq\frac{1}{3}\Longrightarrow\frac{g_{k}(\Delta)}{g_{k+1}(\Delta)}\leq 3$ since we assumed that $g(\Delta)\geq 0$. Substituting this into our inequality above, we have
@@
    This is a stronger bound for $dg\gg 0$.
@@
by Theorem 1.6. This is a necessary condition when $\Delta$ is a simplicial sphere that satisfies the link condition.

### 26 Mar 2026 v3

@@ Introduction
In order to have a better grasp of how we can think about these objects objects quantitatively, we considered averaging and asymptotic properties. This does not seem to have been studied in many settings in the context of gamma vectors apart from work on symmetric edge polytopes in [15]. More specifically, we focused on measuring how important the flag property is to nonnegativity of gamma vectors of simplicial spheres. To this end, we replaced the flag property by the more general link condition $\operatorname{lk}_{\Delta}(a)\cap\operatorname{lk}_{\Delta}(b)=\operatorname{lk}_{\Delta}(e)$ for edges $e\in\Delta$ of a simplicial complex $\Delta$. This is a common combinatorial application of the flag property which ends up being equivalent to preservation of piecewise linear (PL) homeomorphism types of compact PL manifolds without boundary [37]. We will mainly consider PL simplicial spheres.
Given a simplicial sphere, the characterization of possible $g$-vectors of simplicial spheres (Theorem 1.7) and approximations of pseudopowers arising from Macaulay representations (Definition 1.8) discussed in Section 2.1 (e.g. applications of results in [17] and [10]) enable us to study the effect of the link condition on the structure of simplicial spheres with nonnegative gamma vectors in Section 2.2. The results involved also apply to generalizations of gamma positivity coming from connections between orthogonal polynomials and lattice paths (see Section 3). This fits into lattice path-related properties of gamma vector transformations (Proposition 2.3), recent gamma positivity-related work (e.g. [11] and [16]), and positivity and structural results of the gamma polynomial from inverted Chebyshev expansions (Theorem 1.2, Corollary 1.35 on p. 25 of [41] and Corollary 1.10 on p. 9 of [44]). The main results are listed below.
@@
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial sphere satisfying the link condition such that $\gamma(\Delta)\geq 0$. Given an edge $e\in\Delta$, let $\widetilde{\Delta}_{e}$ be the contraction of $e$. Fix $1\leq k\leq\frac{d}{2}-2$ in Part 2 and Part 3.
@@
    For gamma vectors $\gamma(\Delta)$, lower bounds on $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ can be phrased as lower bounds on $\gamma_{k+2}(\Delta)$ since this term only appears in $g_{k+2}(\Delta)$ (see Proposition 2.1). However, the asymptotic effect on high-dimensional spheres satisfying the link condition with $\gamma(\Delta)\geq 0$ is the largest when the gamma vector components involved are asymptotically dominant in the $g$-vector components of the same index. For example, this occurs if $\gamma_{\ell+1}(\Delta)\geq(d+1)g_{\ell}(\Delta)$ (e.g. when $g_{\ell+1}(\Delta)\geq 2(d+1)g_{\ell}(\Delta)$). While the results discussed here involve individual inequalities associated to constant indices $k$, the underlying tools (Proposition 2.40, Corollary 2.42, Corollary 2.45) apply to the behavior of collections of inequalities associated to varying indices.
@@ 1 Background
  3. 3.
    The $g$-vector is defined as
    $g\coloneq(h_{0},h_{1}-h_{0},h_{2}-h_{1},\ldots,h_{\left\lfloor\frac{d}{2}\right\rfloor}-h_{\left\lfloor\frac{d}{2}\right\rfloor-1}).$
@@
Remark 1.6.
(Piecewise linear (PL) assumption)
By “simplicial sphere”, we will mean a piecewise linear (PL) simplicial sphere where the links are spheres (see p. 4 of [27] and p. 26 of [25]).
@@
      The $g$-vector $g\coloneq(h_{0},h_{1}-h_{0},h_{2}-h_{1},\ldots,h_{\left\lfloor\frac{d}{2}\right\rfloor}-h_{\left\lfloor\frac{d}{2}\right\rfloor-1})$ satisfies the inequalities $0\leq g_{i+1}\leq g_{i}^{\langle i\rangle}$ for all $0\leq i\leq\left\lfloor\frac{d}{2}\right\rfloor-1$ (i.e. $g$ is the $h$-vector of some Cohen-Macaulay simplicial complex) (Theorem 2.2 on p. 56 of [46] and Theorem 1.10).
@@
Apart from realizations of simplicial spheres (Theorem 1.7), these expansions are involved in $f$-vectors and $h$-vectors attained by general classes of simplicial complexes. In the first statement, we note that there may a difference in indexing conventions (e.g. including 1-1 index corresponding to index 0 term of an $h$-vector or a gamma vector).
@@ 2.1 General Cohen-Macaulay simplicial complexes and simplicial spheres
In this subsection, we will describe the construction of nonnegative gamma vectors realized by simplicial spheres and Cohen–Macaulay simplicial complexes in general and show that sufficiently large ones can be extended component by component (Corollary 2.13 and Remark 2.14). The key input comes from estimates of pseudopowers from Macaulay representations of $g$-vectors and $h$-vectors (Corollary 2.8) and the $g$-conjecture for simplicial spheres (Theorem 1.7). If $\gamma_{0}=1$ and $\gamma_{1},\ldots,\gamma_{\frac{d}{2}}\geq 0$ that grow sufficiently quickly relative to the dimension with $\gamma_{j}$ while satisfying designated order $\frac{j}{j-1}$ upper bounds in $\gamma_{1},\ldots,\gamma_{j-1}$, they give rise to nonnegative gamma vectors $(\gamma_{0},\gamma_{1},\ldots,\gamma_{\frac{d}{2}})$ realized by simplicial spheres.
@@
    For the first pair of bounds, the term $\frac{a_{r,s}}{a_{r-1,s}}$ can be rewritten as
@@
    where p=d2sp=d-2s and q=rsq=r-s. In our setting, we have $p\geq 2q$ since $0\leq r\leq\frac{d}{2}$. Writing
@@
    $g_{\ell+1}\geq b_{\ell+1,0}+b_{\ell+1,1}\gamma_{1}+\ldots+b_{\ell+1,\ell-1}\gamma_{\ell-1}+b_{\ell+1,\ell}\gamma_{\ell}.$
@@
While there is a lower bound $g_{1}(\Delta)\geq d-1$ for (d1)(d-1)-dimensional simplicial spheres such that $\gamma(\Delta)\geq 0$ (attained by the boundary of the $d$-dimensional cross polytope), the parameter $g_{1}(\Delta)$ can grow arbitrarily quickly with respect to $d$. In other words, $d$ can grow arbitarily slowly with respect to $g_{1}(\Delta)$ even when it is not a constant. We can see this by considering flag simplicial spheres, such that $\gamma(\Delta)\geq 0$ and $\gamma(\operatorname{lk}_{\Delta}(e))\geq 0$ for each edge $e\in\Delta$. This is because flagness is preserved by edge subdivision and each edge subdivision increases the number of vertices by 1.
@@
    This follows from the same reasoning as Part 1 after replacing $g_{i}(\Delta)$ by $h_{i}(\Delta)$, using $a_{i,0}$ instead of $b_{i,0}$ in Proposition 2.1, and applying Theorem 1.10 instead of Theorem 1.7. Note that $h_{i}(\Delta)=1+g_{1}(\Delta)+\ldots+g_{i}(\Delta)$.
@@
Suppose that $\gamma_{k}>0$ for some $1\leq k\leq i$. This would mean that both sides are strictly positive. Let $\alpha=b_{i,0}+b_{i,1}\gamma_{1}+\ldots+b_{i,i-1}\gamma_{i-1}+\gamma_{i}$ and
@@
Taking index changes into account, the $f$-vector realizability conditions of Theorem 1.9 state that we need $\gamma_{i+2}(\Delta)\leq\gamma_{i+1}(\Delta)^{\langle i+1\rangle}$. Substituting in i1i-1 in place of $i$, we have $\gamma_{i+1}(\Delta)\leq\gamma_{i}(\Delta)^{\langle i\rangle}$. The steps of the proof of Corollary 2.13 applying the estimates of Corollary 2.8 imply the controlled growth statement analogous to these results. We will focus on $g$-vectors of simplicial spheres since the analogous statement for $h$-vectors of Cohen–Macaulay simplicial complexes follows from the same reasoning by Lemma 2.3.
@@
If $\Delta$ is a simplicial sphere, Part 2c) of Theorem 1.7 implies that $g(\Delta)=h(T)$ for some Cohen–Macaulay simplicial complex $T$. This implies the following:
@@
    This follows directly from Part 2c) of Theorem 1.7 and Part 3 of Theorem 1.10.
@@ 2.2 Simplicial spheres satisfying the link condition
This is a natural condition to consider from the perspective of piecewise linear (PL) homeomorphisms since they can be decomposed into edge subdivisions and contractions (Corollary [10:2d] on p. 302 of [2], Corollary 4.1 on p. 75 of [31]) and the link condition is equivalent to the contraction of every edge preserving the PL homeomorphism class of a compact PL manifold without boundary [37]. In addition, the condition $\operatorname{lk}_{\Delta}(a)\cap\operatorname{lk}_{\Delta}(b)=\operatorname{lk}_{\Delta}(a,b)$ for edges $\{a,b\}\in\Delta$ is a common application of flagness, which ties many examples from gamma positivity together (see [4], [20]).
@@
In particular, this comes from a shift in perspective on these global/averaged inequalities as information on ratios of $g$-vector components $\frac{g_{\ell+1}(\Delta)}{g_{\ell}(\Delta)}$. Given this situation, we mainly focus on the global/averaged version of the $M$-vector condition on the $g$-vectors of the edge contractions $\widetilde{\Delta}_{e}$ of edges $e\in\Delta$. We note that individual local conditions induce “sandwiching” of local parts (Proposition 2.28) with global/averaged counterparts (Corollary 2.31) when combined with the PL manifold condition. We will primarily focus on the lower bounds and their comparison with nonnegativity of the $g$-vectors of the contractions $\widetilde{\Delta}_{e}$ of edges $e\in\Delta$ (Corollary 2.33) since this is the bound specific to spheres satisfying the link condition rather than PL manifolds in general.
The general shape of the global/averaged $M$-vector conditions on $g$-vectors of contractions $\widetilde{\Delta}_{e}$ of edges $e\in\Delta$ involves a lower bound on $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ in terms of $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$, the number of edges $f_{1}(\Delta)^{\frac{1}{k}}$, and $\frac{g_{k}(\Delta)^{\frac{k+1}{k}}}{g_{k+1}(\Delta)}$ (Proposition 2.38 and Example 2.39). The last parameter measures how far the $M$-vector condition $g_{k+1}(\Delta)\leq g_{k}(\Delta)^{\langle k\rangle}$ is from equality (Part 2c) of Theorem 1.7 and Corollary 2.8). Increasing these parameters decreases the resulting lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$.
@@ 2.2.1 Main underlying inequalities
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial complex satisfying the link condition. Given an edge $e\in\Delta$, let $\widetilde{\Delta}_{e}$ be the contraction of $e$ in $\Delta$. Then, we have
@@
    $\beta_{k}\geq\frac{2\left(2d(d-1)-\left(\frac{d}{2}-1\right)\left(\frac{d}{2}+2\right)\right)}{\frac{d}{2}\left(\frac{d}{2}-1\right)}\approx 14.$
@@
    Suppose that $\Delta$ is a simplicial sphere satisfying the link condition such that $\gamma(\Delta)\geq 0$. In order for the “averaged” global version of the inequalities $g(\widetilde{\Delta}_{e})\geq 0$ for all the edges $e\in\Delta$ from Corollary 2.25 to impose a nontrivial condition on the $g_{k+2}(\Delta)$ beyond the $M$-vector condition on the $g$-vectors of arbitrary simplicial spheres (Part 2c) of Theorem 1.7), we need $g_{2}(\Delta)<(d^{2}-2)-(d-3)g_{1}(\Delta)$. By the $g(\Delta)\geq 0$ assumption, this also implies that $g_{1}(\Delta)<(d+3)+\frac{7}{d-3}$. If $d\geq 4$, the latter bound implies that $g_{1}(\Delta)<d+10$. This implies that $g_{2}(\Delta)\leq\binom{d+11}{2}$, which is a stronger bound for $d\gg 0$. Since there is a linear bound on $g_{1}(\Delta)$, each $g_{\ell}(\Delta)$ with $0\leq\ell\leq\frac{d}{2}$ is bounded above by a degree $\ell$ polynomial in $d$ (Corollary 2.11).
@@
    Since the coefficient of each $g$-vector component in Part 1 of Corollary 2.25 is strictly positive and $g(\Delta)\geq 0$, having $g_{k+1}(\Delta)=0$ would imply that $g_{k+2}(\Delta)=0$ and $g_{k}(\Delta)=0$. If $g_{\ell}(\Delta)=0$ for some $1\leq\ell\leq\frac{d}{2}$, this would mean that $g_{k}(\Delta)=0$ for all $0\leq k\leq\frac{d}{2}$. However, this is impossible in our setting.
@@
    We can compare these to initial bounds from the $M$-vector condition on $g$-vectors of arbitrary simplicial spheres. Since $g_{k+2}(\Delta)\leq g_{k+1}(\Delta)^{\langle k+1\rangle}\leq g_{k+1}(\Delta)^{\frac{k+2}{k+1}}$ and $g_{k+1}(\Delta)\leq g_{1}(\Delta)^{k+1}\Longrightarrow g_{k+1}(\Delta)^{\frac{1}{k+1}}\leq g_{1}(\Delta)$ by Corollary 2.8 and Corollary 2.9, we have
@@
    If $\gamma(\Delta)\geq 0$, then Corollary 2.4 implies that $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}\geq\frac{1}{3}\Longrightarrow\frac{g_{k}(\Delta)}{g_{k+1}(\Delta)}\leq 3$ since we assumed that $g(\Delta)\geq 0$. Substituting this into our inequality above, we have
@@
    This is a stronger bound for $d\gg 0$.
@@
by Theorem 1.7. This is a necessary condition when $\Delta$ is a simplicial sphere (Remark 1.6) that satisfies the link condition.